# A way around the exponential scaling in optimal quantum control

### v1

@@ Abstract
We show that combining ideas from the fields of quantum invariants and of optimal control can be used to design quantum control of quantum systems without explicit reference to quantum states. The scaling in numerical effort of the resultant approach is given by commutation relations of system operators, and it can be polynomial in the number of subsystems despite the general quantum mechanical exponential scaling of the Hilbert space. As explicit applications, we discuss state preparation and quantum simulation with Hamiltonians including three-body and many-body interactions with spin chains of up to $50$ constituents, and the perspective of use for topologically protected quantum information processing.
@@ I Introduction
The availability of well-controlled quantum systems offers the prospect towards a broad range of applications, including quantum simulation [1, 2] and quantum computation [3, 4]. The exponential growth of the Hilbert space in such systems suggests that even moderately small quantum systems could be employed to achieve tasks that would be impractical to perform classically [5, 6]. While the benefits of this scaling have inspired tremendous efforts towards the development of quantum technological devices [7, 8, 9, 5], the scaling also has disastrous implications on our abilities to simulate the dynamics of quantum systems [10].
@@
A very efficient, even though implicit way of describing some quantum states is given in terms of the operator (or operators) to which a state is an eigenstate [20]. There is, for example, a large variety of quantum many-body Hamiltonians with highly complicated ground states [21, 22, 23]. The Hamiltonian is typically specified very efficiently in terms of the interaction geometry and the type of interaction mechanism. Even though the explicit specification of the ground state of such a Hamiltonian is hardly ever efficient, the implicit specification via the Hamiltonian is.
While ideas to characterize quantum states in terms of operators to which they are eigenstates do exist [24], optimal control seems to exclusively focus on explicit descriptions of state vectors [25]. The goal of this paper is to demonstrate that optimal control can be realized without explicit construction of quantum states, and that this can make control problems numerically accessible whose realization based on explicit state vector descriptions are prohibitively expensive.
The underlying idea is inspired by the framework of quantum invariants [26, 27, 28], i.e. hermitian operators ${\cal I}(t)$ with a vanishing total time-derivative, i.e.
$\frac{d{\cal I}}{dt}=\frac{\partial{\cal I}}{\partial t}+i[H(t),{\cal I}(t)]=0\ .$ (1)
@@
$\mathcal{I}(t)\ket{\Psi(t)}=\lambda\ket{\Psi(t)}\ .$ (2)
Due to the unitary dynamics resultant from Eq. (1), any eigenvalue $\lambda$ of $\mathcal{I}(t)$ is time-independent.
Since every non-degenerate, instantaneous eigenvector specifies a solution of the time-dependent Schrödinger equation with the time-dependent Hamiltonian $H(t)$ [26], invariants are commonly used tools in optimal quantum control [29].
Existing use of quantum invariants in optimal quantum control is based on the identification of analytic time-dependences that make an operator satisfy Eq. (1) for a certain type of Hamiltonian, such as the driven harmonic oscillator Hamiltonian and the Hamiltonian of a single driven spin [30, 31, 32]. Once such an analytic form is identified, the resultant invariant provides very elegant means of engineering a time-dependent Hamiltonian that induces desired dynamics. It does, however, seem elusive to generalize this analytic approach to more complicated systems such as interacting spins.
Our present goal is thus to develop a framework in the spirit of numerical quantum control [4, 33] that does not rely on the existence of analytic solutions, but that avoids reference to quantum states whose description would require an exponential effort.
@@ II Implicit quantum control
The goal of this manuscript is to show how to define a control target ${\cal I}_{T}$ for an operator ${\cal I}(t)$ following the dynamics of Eq. (1), such that numerical pulse shaping techniques can be applied to Eq. (1), and such that the resultant Hamiltonian induces desired dynamics in the underlying Schrödinger equation.
@@ II.1 Operator expansion
The crucial aspect to determine whether the present approach is more efficient than approaches based on state vectors is the effort required to perform the propagation of ${\cal I}(t)$ following Eq. (1).
@@
$H(t)=\sum_{j}h_{j}{\mathfrak{h}}_{j}\ ,$ (3)
spanned by a set of operators ${\mathfrak{h}}_{j}$ with scalar expansion coefficients $h_{j}$. Some of those can be time-independent and fixed (as in a drift Hamiltonian), but some should be time-dependent and tuneable (as in a control Hamiltonian).
The operator
${\cal I}=\sum_{j=1}^{d}a_{j}{\mathfrak{a}}_{j}\ ,$ (4)
to be propagated, is expanded into a set of time-independent operators ${\mathfrak{a}}_{j}$ with time-dependent scalar expansion coefficients $a_{j}$. This Ansatz can help to solve the equation of motion Eq.(1), only if the set of operators $\{{\mathfrak{a}}_{j}\}$ is closed under commutation with any of the operators ${\mathfrak{h}}_{k}$, i.e. if any commutator $[{\mathfrak{a}}_{j},{\mathfrak{h}}_{k}]$ can be expanded in terms of the set $\{{\mathfrak{a}}_{j}\}$, such that
$[{\mathfrak{a}}_{j},{\mathfrak{h}}_{k}]=-i\sum_{l=1}^{d}\lambda_{jl}^{k}{\mathfrak{a}}_{l}\ .$ (5)
In any finite-dimensional system, this condition can always be satisfied, if the set of operators is chosen to have sufficiently many elements, but in the following, it will be crucial that the set $\{{\mathfrak{a}}_{j}\}$ is sufficiently small.
If the sets of operators $\{{\mathfrak{a}}_{j}\}$ and $\{{\mathfrak{h}}_{k}\}$ coincide, the present discussion reduces to the common structure of a Lie algebra, but in order to find small sets of operators $\{{\mathfrak{a}}_{j}\}$ it can be beneficial to consider situations beyond Lie algebras.
II.2 Equations of motion
@@
$\frac{\partial{\cal I}}{\partial t}=i\sum_{j,k}a_{j}h_{k}\commutator{\mathfrak{a}_{j}}{\mathfrak{h}_{k}}=\sum_{j,k,l}a_{j}h_{k}\lambda_{jl}^{k}\mathfrak{a}_{l}\ .$ (6)
@@
II.3 Formulation of the control problem
@@
While typically control problems are defined in terms of a target state or a target gate, the present framework requires a control target for the operator ${\cal I}$, and defining this requires some care. In the following, we will discuss this for the problems of diabatic transitions between eigenstates of two operators and for the realization of a desired unitary.
II.3.1 Diabatic transitions
The problem of diabatic transitions is a natural application of the present framework, because Hamiltonians can typically be specified efficiently, whereas the explicit specification of their eigenstates is often not efficient.
For the problem of finding a time-dependent Hamiltonian such that the system evolves from the ground-state of an initial operator $H_{I}$ to the ground-state of a final operator $H_{F}$, it seems tempting to define the initial condition ${\cal I}(0)=H_{I}$ and the control target ${\cal I}_{T}=H_{F}$. However, since the spectrum of ${\cal I}(t)$ is time-independent, this would necessarily imply the restriction that $H_{I}$ and $H_{F}$ have the same spectrum.
In addition to the natural requirements that
@@ II.3.1 Diabatic transitions
    the initial condition ${\cal I}(0)$ needs to be specified such that the initial state |$\ket{\Psi_{1}}$ (an eigenvector of $H_{I}$) is an eigenstate of ${\cal I}(0)$ with a non-degenerate eigenvalue $\gamma$; and that
@@
    the control target ${\cal I}_{T}$ needs to be chosen such that the target state |$\ket{\Psi_{2}}$ (an eigenvector of $H_{F}$) is an eigenstate of ${\cal I}_{T}$ with the same eigenvalue $\gamma$,
there is thus the third requirement that
@@
    the control target needs to be reachable with the dynamics induced by the control Hamiltonian $H(t)$, i.e., the dynamics following Eq. (8).
If the choice ${\cal I}(0)=H_{I}$ is taken, one can consider an adiabatic transition from $H_{I}$ to $H_{F}$, and simulating the adiabatic evolution in Eq. (8) yields a final invariant that is reachable with the available dynamics and that has |$\ket{\Psi_{2}}$ as eigenstate. While such a numerical construction is a generally applicable approach, there are also instances in which suitable control targets can be defined analytically as exemplified in Secs. III.2.1 and III.2.2.
@@
II.3.2 Desired unitary
In contrast to the above problem of diabatic transitions, there is no ambiguity in defining the target operator ${\cal I}_{T}$ corresponding to an initial condition ${\cal I}(0)$ for the control problem of realizing a desired unitary ${\cal U}_{T}$. The problem is that optimizing for an objective function ${\cal I}_{T}={\cal U}_{T}{\cal I}(0){\cal U}_{T}^{\dagger}$ does not ensure that the desired unitary is actually achieved.
An objective function such that its optimal value unambiguously verifies the realization of the desired unitary ${\cal U}_{T}$ requires a set of $N_{c}$ initial conditions ${\cal I}_{j}(0)$ and corresponding targets $\mathcal{I}_{T,j}={\cal U}_{T}{\cal I}_{j}(0){\cal U}_{T}^{\dagger}$. However, only with a sufficiently large number of these initial conditions does the objective function,
guarantee that a vanishing value of ${\cal J}_{{\cal U}_{T}}$ unambiguously implies the realized unitary coincides with the desired unitary ${\cal U}_{T}$.
While, in general, the gate-infidelity ${\cal J}_{{\cal U}_{T}}$ would require an exponentially large set of initial conditions, the fact that the dynamics takes place in a Lie group of small size also permits to define a faithful target that requires a correspondingly small number of initial conditions, as exemplified further down in Sec. IV.1. With each summand in Eq. (14) evaluated analogously to Eq. (12), the objective function ${\cal J}_{{\cal U}_{T}}$ can thus be evaluated with polynomial effort.
II.3.3 Numerical optimization
Once a control target ${\cal I}_{T}$ and an objective function, such as ${\cal J}$ (Eq. (12)) or ${\cal J}_{{\cal U}_{T}}$ (Eq. (14)) is specified, one can use any numerical pulse shaping algorithm [35] to design a time-dependent Hamiltonian that realizes a dynamics from ${\cal I}(0)$ to ${\cal I}(T)$, at a final time $T$, that is close to ${\cal I}_{T}$. Since Eq. (8) with an anti-symmetric matrix $K$ induces orthogonal (unitary and real) dynamics, any numerical algorithm that is applicable to the Schrödinger equation (such as Krotov [36], GRAPE [11] or CRAB [14]) is directly applicable, and techniques to construct e.g. gradients with respect to control parameters also directly apply to Eq. (8).
Mere continuity implies that in the limit of vanishing infidelity ${\cal J}$ for the non-degenerate operator ${\cal I}(T)$ also the infidelity of the realized quantum state approaches its ideal value. Beyond this qualitative assessment, also a more quantitative estimate can be obtained in terms of the spectrum of $H_{F}$.
To this end, the final state |$\ket{\psi(T)}$ can be expressed as
@@ II.3.3 Numerical optimization
with the target state |$\ket{\psi_{T}}$, i.e. the ground state of $H_{F}$, and the projector $P_{e}$ onto the space spanned by the excited states of $H_{F}$.
The expectation value of $H_{F}$ with respect to |$\ket{\psi(T)}$ is
$\bra{\psi(T)}H_{F}\ket{\psi(T)}=|\innerproduct{\psi_{T}}{\psi(T)}|^{2}E_{0}+\bra{\phi}H_{F}\ket{\phi}\ ,$ (16)
with |$\ket{\phi}=P_{e}\ket{\psi(T)}$ and the ground state energy $E_{0}$ of $H_{F}$. The norm of the vector |$\ket{\phi}$ is given in terms of the state fidelity $F=|\innerproduct{\psi_{T}}{\psi(T)}|^{2}$ via the relation
@@
in terms of the energy $E_{1}$ of the first excited state of $H_{F}$, the state fidelity is bounded by
$1-F\leq\frac{\bra{\psi(T)}H_{F}\ket{\psi(T)}-E_{0}}{E_{1}-E_{0}}\ .$ (19)

### v2

# Quantum control without quantum states

@@ Abstract
We show that combining ideas from the fields of quantum invariants and of optimal control can be used to design optimal quantum control solutions without explicit reference to quantum states. The states are specified only implicitly in terms of operators to which they are eigenstates. The scaling in numerical effort of the resultant approach is not given by the typically exponentially growing effort required for the specification of a time-evolved quantum state, but it is given by the effort required for the specification of a time-evolved operator. For certain Hamiltonians, this effort can be polynomial in the system size.
We describe how control problems for state preparation and the realization of propagators can be formulated in this approach, and we provide explicit control solutions for a spin chain with an extended Ising Hamiltonian. The states considered for state-preparation protocols include eigenstates of Hamiltonians with more than pairwise interactions, and these Hamiltonians are also used for the definition of target propagators. The cost of describing suitable time-evolving operators grows only quadratically with the system size, allowing us to construct explicit control solutions for up to $50$ spins. While sub-exponential scaling is obtained only in special cases, we provide several examples that demonstrate favourable scaling beyond the extended Ising model.
@@ I Introduction
The availability of well-controlled quantum systems offers the prospect towards a broad range of applications [1], including quantum simulation [2, 3] and quantum computation [4, 5, 6]. The exponential growth of the Hilbert space in such systems suggests that even moderately small quantum systems could be employed to achieve tasks that would be impractical to perform classically [4, 7]. While the benefits of this scaling have inspired tremendous efforts towards the development of quantum technological devices [6, 8, 9, 10, 11], the scaling also has disastrous implications on our abilities to simulate the dynamics of quantum systems [12].
@@
A very efficient, even though implicit way of describing some quantum states is given in terms of the operator (or operators) to which a state is an eigenstate [22]. There is, for example, a large variety of quantum many-body Hamiltonians with highly complicated ground states [23, 24, 25]. The Hamiltonian itself is specified in terms of the interaction geometry, the type of interaction mechanism and some parameters specifying quantities like interaction strengths. For many systems, the number of parameters required for this specification scales polynomially with the number of interacting objects; the description is thus efficient. The ground state of many such Hamiltonians, however, often lacks an efficient representation, i.e. a representation specified with a polynomial number of parameters. Implicitly, however, the ground state is specified via the Hamiltonian. This implicit description of states is foundational to the stabilizer formalism [26] and to adiabatic quantum computation [27].
Building on this concept of implicit state characterization, this paper introduces a framework for optimal control that avoids explicit construction of quantum states, while not being limited to adiabatic time scales. Formulating an optimal quantum control formalism in terms of operators, we demonstrate a method that can help solving control problems that are computationally intractable using traditional state-vector-based approaches, and that does not require an approximate treatment.
The underlying idea is inspired by the framework of quantum invariants [28, 29, 30], i.e. hermitian operators ${\cal I}(t)$ satisfying the equation of motion
$\frac{\partial{\cal I}}{\partial t}+i[H(t),{\cal I}(t)]=0\ ,$ (1)
in the Schrödinger picture.
@@
$\mathcal{I}(t)\ket{\Psi(t)}=\gamma\ket{\Psi(t)}\ .$ (2)
Due to the unitary dynamics resultant from Eq. (1), any eigenvalue $\gamma$ of $\mathcal{I}(t)$ is time-independent. Since every non-degenerate, instantaneous eigenvector specifies a solution of the time-dependent Schrödinger equation with the time-dependent Hamiltonian $H(t)$ [28], invariants are commonly used tools in optimal quantum control [31].
Existing use of quantum invariants in optimal quantum control is based on the identification of analytic time-dependences that make an operator satisfy Eq. (1) for a certain type of Hamiltonian, such as the driven harmonic oscillator Hamiltonian and the Hamiltonian of a single driven spin [32, 33, 34]. Once such an analytic form is identified, the resultant invariant provides very elegant means of engineering a time-dependent Hamiltonian that induces desired dynamics.
Finding such analytic expressions is practical only for particularly simple cases, and it does seem elusive to find an analytic invariant for quantum systems with a large number of interacting subsystems. The goal of this paper is thus to develop an approach that benefits from some aspects of quantum invariants, but that does not rely on the existence of an analytic form of an invariant. This approach will combine elements from optimal control based on quantum invariants [31] and from numerical quantum control with state vectors or propagators [1, 35]. It relies on iterative numerical refinement of a time-dependent Hamiltonian, but instead of the commonly employed numerical integration of the Schrödinger equation, the present approach is based on numerical integration of the equation of motion of quantum invariants, i.e. Eq. (1). Since such numerical integration does not require the existence of analytic solutions, this approach can indeed be applicable to interacting quantum systems.
Instead of defining a control target for state vectors or propagators, control targets are defined for operators ${\cal I}(t)$, satisfying the equation of motion of quantum invariants Eq. (1). This avoids reference to quantum states with corresponding exponential effort, and, as shown with explicit examples, there are interacting quantum systems for which the description of the time-dependent operator $\mathcal{I}(t)$ does not require an exponential effort.
@@ II Implicit quantum control
In this section we show how to define a control target ${\cal I}_{T}$ for an operator ${\cal I}(t)$ following the dynamics of Eq. (1), such that numerical pulse shaping techniques can be applied to Eq. (1), and such that the resultant Hamiltonian induces desired dynamics in the underlying Schrödinger equation.
@@ II.1 Operator expansion
The crucial aspect to determine whether the present approach is more efficient than approaches based on state vectors is the dimension of the space in which the dynamics of ${\cal I}(t)$ following Eq. (1) takes place.
@@
$H(t)=\sum_{j=1}^{d_{0}}h_{j}(t){\mathfrak{h}}_{j}\ ,$ (3)
spanned by a fixed set of $d_{0}$ operators ${\mathfrak{h}}_{j}$ with scalar expansion coefficients $h_{j}(t)$. Some of those can be time-independent and fixed (as in a drift Hamiltonian), but some should be time-dependent and tuneable (as in a control Hamiltonian).
The Ansatz
${\cal I}(t)=\sum_{j=1}^{d}a_{j}(t){\mathfrak{a}}_{j}$ (4)
for the operator ${\cal I}(t)$ to be propagated with a set of $d$ time-independent operators ${\mathfrak{a}}_{j}$ and time-dependent scalar expansion coefficients $a_{j}(t)$ is sufficient to solve Eq. (1) for any choice of the time-dependent coefficients $h_{j}(t)$, if the commutation relations
$[{\mathfrak{a}}_{j},{\mathfrak{h}}_{k}]=-i\sum_{l=1}^{d}\lambda_{lj}^{k}{\mathfrak{a}}_{l}\ ,$ (5)
with real coefficients $\lambda_{lj}^{k}$ are satsfied for all $k\in[1,d_{0}]$ and $j\in[1,d]$.
An Ansatz given by Eq. (4) that does not lead to the closed commutator relations of Eq. (5) can be salvaged with an increase of the set of operators ${\mathfrak{a}}_{j}$, i.e. and increase of $d$. Given a set of $d$ operators ${\mathfrak{a}}_{j}$ with a pair of operators ${\mathfrak{a}}_{j_{0}}$ and ${\mathfrak{h}}_{k_{0}}$ such that the commutator $[{\mathfrak{a}}_{j_{0}},{\mathfrak{h}}_{k_{0}}]$ can not be expanded in terms of the set of $d$ operators ${\mathfrak{a}}_{j}$, one can define ${\mathfrak{a}}_{d+1}=i[{\mathfrak{a}}_{j_{0}},{\mathfrak{h}}_{k_{0}}]$ as additional element, increasing the value of $d$ by $1$. In any finite-dimensional system, this process can be continued until Eq. (5) is satisfied for all pairs $\mathfrak{a}_{j}$, $\mathfrak{h}_{k}$.
In practice, this construction can be initiated with a single element $\mathfrak{a}_{1}$, i.e. with $d=1$, and the eventual value of $d$ depends on the choice of initialization. Since the effort to integrate Eq. (1) grows with increasing $d$, it is desirable to find small sets of operators $\mathfrak{a}_{j}$ with closed commutator relationships, but even with suitable choices, $d$ will typically be larger than $d_{0}$.
II.2 Lie algebra
The construction of the set of operators $\{\mathfrak{a}_{j}\}$ discussed above is a slight variant of the construction of a Lie-algebra, i.e. a set of $d$ operators $\mathfrak{h}_{j}$ with a closed commutator relationship $[{\mathfrak{h}}_{j},{\mathfrak{h}}_{k}]=-i\sum_{l=1}^{d}\lambda_{lj}^{k}{\mathfrak{h}}_{l}$ for all $j,k\in[1,d]$. Similarly to the construction above, it can be initiated with a set of $d_{0}$ operators with $d_{0}<d$ and the construction is completed when all commutation relations are closed.
Crucially, the propagator induced by the time-dependent Hamiltonian $H(t)$ (Eq.(3)) is not necessarily of the form $U(t)=\exp(-i\sum_{j=1}^{d_{0}}b_{j}(t)\mathfrak{h}_{j})$ with time-dependent scalar functions $b_{j}(t)$ and the terms $\mathfrak{h}_{j}$ ($j\in[1,d_{0}]$) in the system Hamiltonian, but typically also terms with $j>d_{0}$ of the Lie algebra contribute. In turn, a propagator of the form $U=\exp(-iK)$ with an operator $K$ that cannot be expanded in terms of the basis elements of the Lie algebra (i.e. $\mathfrak{h}_{j}$, $j\in[1,d]$) cannot be realized with the dynamics induced by $H(t)$ in Eq. (3) and is thus considered not reachable [36].
In the explicit examples Sec. III and IV the time-evolving operator will be expanded in terms of the Lie algebra generated by the operators $\mathfrak{h_{j}}$ with $j\in[1,d_{0}]$. In order to find small sets of operators $\{{\mathfrak{a}}_{j}\}$ it can, however, also be beneficial to consider situations beyond Lie algebras, as exemplified in Sec. V.
II.3 Equations of motion
@@
$\frac{\partial{\cal I}}{\partial t}=i\sum_{j,k}a_{j}(t)h_{k}(t)\commutator{\mathfrak{a}_{j}}{\mathfrak{h}_{k}}=\sum_{j,k,l}a_{j}(t)h_{k}(t)\lambda_{lj}^{k}\mathfrak{a}_{l}\ .$ (6)
@@
II.4 Formulation of the control problem
@@
While typically control problems are defined in terms of a target state or a target gate, the present framework requires a control target for the operator ${\cal I}$, and defining this requires some care. In the following, we will discuss this for the problems of state preparation and for the realization of propagators.
II.4.1 State preparation
The problem of state preparation is a natural application of the present framework, because Hamiltonians can typically be specified efficiently, whereas the explicit specification of their eigenstates is often not efficient.
With a control problem defined in terms of a target state, the target functional is naturally defined in terms of this state. The present approach leaves much more freedom, since there are many operators ${\cal I}_{T}$ that have a given target state as ground state. There are thus many conceivable choices for ${\cal I}_{T}$ and a good choice is a central part of the control problem within the present approach.
For the problem of finding a time-dependent Hamiltonian such that the system evolves from the ground-state of an initial operator $H_{I}$ to the ground-state of a final operator $H_{T}$, it seems tempting to define the initial condition ${\cal I}(0)=H_{I}$ and the control target ${\cal I}_{T}=H_{T}$. However, since the spectrum of ${\cal I}(t)$ is time-independent, this would necessarily imply the restriction that $H_{I}$ and $H_{T}$ have the same spectrum.
Moreover, if the operators $\mathfrak{h}_{j}$ in the system Hamiltonian (Eq.(3)) do not generate the maximal Lie algebra, i.e. an algebra that contains a complete set of operators apart from the identity, then there are unitary propagators that cannot be realised with the system Hamiltonian, as discussed in the penultimate paragraph of Sec. II.2. That means that even if $H_{I}$ and $H_{T}$ have the same spectrum, it is not necessarily possible to realise the dynamics such that an operator ${\cal I}(t)$ with the initial condition ${\cal I}(0)=H_{I}$ evolves towards the final value ${\cal I}(T)=H_{T}$; i.e. the target ${\cal I}_{T}=H_{T}$ is not reachable (explicit examples demonstrating that specific targets can be reached are discussed in Secs. III.2.1 and III.2.2).
It is thus necessary to impose the requirements that
@@ II.4.1 State preparation
    the initial condition ${\cal I}(0)$ needs to be specified such that the initial state |$\ket{\Psi_{I}}$ (an eigenvector of $H_{I}$) is an eigenstate of ${\cal I}(0)$ with a non-degenerate eigenvalue $\gamma$; and that
@@
    the control target ${\cal I}_{T}$ needs to be specified such that the target state |$\ket{\Psi_{T}}$ (an eigenvector of $H_{T}$) is an eigenstate of ${\cal I}_{T}$ with the same eigenvalue $\gamma$,
@@
    the control target needs to be reachable with the dynamics induced by the system Hamiltonian $H(t)$, i.e., the dynamics following Eq. (8).
In attempting to satisfy these requirements it can be helpful to consider the following aspects:
For (i): In most control problems, the initial operator $H_{I}$ would be a non-interacting operator, i.e. a sum of single-qubit terms (or, more generally single-particle terms) without any interactions. As such, also the initial condition ${\cal I}(0)$ can be chosen to be non-interacting. As long as the single-qubit parts of ${\cal I}(0)$ are non-degenerate, such an operator has a unique, non-degenerate ground-state, and the associated eigenvalue is given by the sum of all the single-qubit ground-state eigenvalues. In cases in which the initial operator ${\cal I}(0)$ is interacting, it might be challenging to verify the non-degeneracy of the ground state. There is certainly no unique recipe to do this, but one option would be to show unitary equivalence of this operator with a non-degenerate, non-interacting operator.
For (ii): There is a general approach to numerically construct a reachable control target that is applicable if the set of operators $\{\mathfrak{a}_{j}\}$ is taken to coincide with the Lie algebra generated by the operators $\mathfrak{h}_{j}$ defining the system Hamiltonian in Eq. (3). Given any initial condition ${\cal I}(0)$, one can numerically simulate the dynamics of ${\cal I}(t)$ under a Hamiltonian that evolves adiabatically from $H_{I}$ to $H_{T}$, as demonstrated in Sec. III.2.3. In practice, this implies numerical integration of Eq. (8), with an effort that scales with $d$ and not with the exponentially large dimension of the system Hilbert space. The final operator ${\cal I}_{T}$ obtained in this fashion necessarily has the same spectrum as ${\cal I}(0)$ because of unitarity of the dynamics, its ground state coincides with the ground state of $H_{T}$ because of adiabaticity, and it is reachable with the dynamics induced by the system Hamiltonian.
For (iii): Given Eq. (8), the question of reachability of an operator ${\cal I}_{T}$ starting from an operator ${\cal I}(0)$ is equivalent to the question of reachability of a vector ${\bf a}^{T}$ (such that ${\cal I}_{T}=\sum_{j}a_{j}^{T}\mathfrak{a}_{j}$) starting from a vector ${\bf a}(0)$ (such that ${\cal I}(0)=\sum_{j}a_{j}(0)\mathfrak{a}_{j}$) given the time-dependent operators $K(t)$ (Eq. (9)). The question of reachability at hand is thus equivalent to standard reachability of vectors with a linear differential equation, as in the case of state control with the Schrödinger equation [38].
The requirements (i), (ii) and (iii) do not result in a unique choice of initial condition ${\cal I}(0)$ and target ${\cal I}_{T}$, and this freedom of choice can be used to select an initial condition ${\cal I}(0)$ that can be expanded in a set of operators $\{\mathfrak{a}_{j}\}$ such that the commutation relations Eq. (5) are closed even though $d$ is small.
Some of the examples discussed in Sec. III, will rely on the numerical construction of control targets with simulated adiabatic dynamics. While such a numerical construction is a generally applicable approach, there are also instances in which suitable control targets can be defined analytically as exemplified in Secs. III.2.1 and III.2.2.
II.4.2 State fidelity
@@
If ${\cal I}(T)$ coincides exactly with the control target ${\cal I}_{T}$, then it is ensured that the system state coincides with the desired non-degenerate eigenstate of ${\cal I}_{T}$, i.e. with the target state. Since both state infidelity and the infidelity defined in Eq. (12) are continuous functions, continuity implies that in the limit of vanishing infidelity ${\cal J}$ for the non-degenerate operator ${\cal I}(T)$ also the infidelity of the realized quantum state approaches its ideal value. In practice, a more quantitative estimate is desirable to assess the quality of an optimal control solution.
In order to obtain such an estimate that will depend on the spectrum of $H_{T}$, it is helpful to express the final state |$\ket{\Psi(T)}$ as
@@
with the target state |$\ket{\Psi_{T}}$, i.e. the ground state of $H_{T}$, and the projector $P_{e}$ onto the space spanned by the excited states of $H_{T}$.
The expectation value of $H_{T}$ with respect to |$\ket{\Psi(T)}$ reads
$\bra{\Psi(T)}H_{T}\ket{\Psi(T)}=|\innerproduct{\Psi_{T}}{\Psi(T)}|^{2}E_{0}+\bra{\phi}H_{T}\ket{\phi}\ ,$ (15)
with |$\ket{\phi}=P_{e}\ket{\Psi(T)}$ and the ground state energy $E_{0}$ of $H_{T}$. The norm of the vector |$\ket{\phi}$ is given in terms of the state fidelity $F=|\innerproduct{\Psi_{T}}{\Psi(T)}|^{2}$ via the relation
@@
in terms of the energy $E_{1}$ of the first excited state of $H_{T}$, the state fidelity is bounded by
$1-F\leq\frac{\bra{\Psi(T)}H_{T}\ket{\Psi(T)}-E_{0}}{E_{1}-E_{0}}\ .$ (18)